A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.00 s for the ball to reach its maximum height. Find (a) the ball’s initial velocity and (b) the height it reaches.
1
Expert's answer
2020-10-07T10:23:46-0400

(a)

"v=gt=9.8\cdot3=29.4" ("\frac{m}{s}" )

(b)

"h=\frac{gt^2}{2}=\frac{9.8\cdot3^2}{2}=44.1" ("m" )